# Instructions for *ACL Proceedings

### v1

@@ 1 Introduction
Recently, Large Language Models (LLMs) have become the leading architecture for many tasks in Natural Language Processing (NLP) Kojima et al. (2022); Zhou et al. (2022); Chang et al. (2024). Despite their class-leading performance, LLMs have been shown to propagate and amplify different forms of bias in numerous domains Wan et al. (2023); Gupta et al. (2023); Poulain et al. (2024), similar to how more traditional predictive machine learning-based models replicate and at times exacerbate social biases Mehrabi et al. (2021).
@@
    Candidates with White names are preferred by GPT-3.5-Turbo and Llama 3 over other groups in between 50% to 95% of 40 occupations, depending on the setting and model.
@@
    Even when candidates posses identical qualifications as reflected in biographies, the average salary offered by these LLMs to candidates with female names may still differ up to 1.8% compared to male counterparts. This discrepancy reaches up to 5% when comparing candidates from intersectional groups.
@@
    Biases exhibited by LLMs partially mirror real-world trends in the United States (U.S) labor force at coarse-grain levels. However, intersectional analysis reveals nuanced discrepancies that favor certain minority groups while punishing others.
Our work builds directly on that of Haim et al. (2024) and of An et al. (2024). Haim et al. (2024) prompted LLMs to provide assistance for 40 Black and White named individuals across topics related to sports, public office, purchasing etc., finding that Black female names received the worst outcomes. An et al. (2024) prompted LLMs to write emails to accept or reject job candidates with stereotypically White, Black or Hispanic names (across two genders), and investigated whether those emails chose to accept or reject the candidates. Their work found that acceptance rates for the latter 2 groups tend to be lower than the former, even when degrees of education and qualification level were consistently stated across candidates. Our work augments these findings by: 1) exploring alternative hiring-related tasks, including salary prediction with full, natural biographies—similar to “résumé studies” in sociology—, and 2) by connecting LLM behaviors to real-world labor data to reveal intersectional bias by occupation.
@@ 2 Hiring Recommendation
In this paper, we study two types of recommendations that LLMs could conceivably be used for. The first, discussed in this section, is hiring recommendations: given an occupation and a list of names of potential candidates, do LLMs exhibit any racial or gender preferences for selecting who to give a job to ? The second type is salary recommendation (§ 3): given a candidate name and (potentially) a biography for that candidate, what salary is recommended for them ?
@@ 2.1 Curation of Names
We limit our study to first (given) names, drawn from two datasets: Rosenman et al. (2023) and Tzioumis (2018). The former contains 136,000 first names compiled form voter registration files while the latter contains 4,250 first names compiled mortgage information in the United States. Both datasets supply the conditional probabilities of each name’s association with 5 racial/ethnic groups: White, Black, Hispanic, Asian, Other. We designate a name as male or female based on its probability in the U.S. Social Security Agency’s database, with a 0.5 probability threshold. Omitting the Other category due to its small size, we randomly select 40 names from each gender of the four remaining races, whose conditional probability P(race|name) exceeds at least 0.8. Our final pool consists of 320 first names. See § A.1 for additional details.
@@ 2.3 Hiring Recommendation Results
First, LLMs show a strong preference for White names, particularly favoring White female names over other groups. For gender-stratified hiring, White female names are preferred in more occupations (35 by GPT-3.5, 29 by Llama 3) compared to White male names (30 and 18) (Table 1). For gender-inclusive hiring, White female names are preferred in 28 (70%) and 26 (65%) occupations by GPT-3.5 and Llama 3 (Table 2)
Second, Llama 3 exhibits less bias for White names compared to GPT-3.5. In Table 1, Asian names are the second most chosen group across occupations, though not significantly so. In contrast, Black names are disproportionately hired as rapper by GPT-3.5, with the addition of singer and social worker by Llama 3. Hispanic names are never the majority for any occupation by GPT-3.5, and only for 5 and 2 occupations among male and female groups by Llama 3. In Table 2, Llama 3 exhibits more distributed preference for non-White names vs. GPT-3.5, though still far from parity.
@@ 3.2.2 Intersectional Analysis
We see that LLM-projected gender salary gaps are still significantly less than U.S data’s on average. The 2022 ACS reports that females make more than males in only 3 of 16 occupations (dietitian, interior designer, paralegal), with the average absolute percentage gap between the median salaries of the 2 genders at 13.03% (Table 9). In contrast, the average gender gaps between LLMs’ recommended mean salaries are all less than 1.01 $\pm$ 0.82% (Table 6). Interestingly, the average MAEs with respect to U.S statistics remain consistent around 12 units for both LLMs, with their variance also similar.
@@ 4 Discussion
For instance, female names are preferred over male names in gender-inclusive hiring, yet often offered less salary for the same position than their male counterparts. In contrast, Black names are often overlooked in hiring, but are also offered salary higher than average. In comparison, White names are consistently preferred in both hiring and salary recommendation, with Hispanic names often on the opposite end.
@@
Our results showcase that prompting LLMs to choose one among several candidates arguably magnify the risk of preferential treatment, and thus should be avoided. Though Llama 3 displays larger magnitude of bias than GPT-3.5, its open-source nature lends itself to more mitigation strategies (more detailed discussion in Appendix A.5 ) Zhou et al. (2023); Qureshi et al. (2023); Wang and Russakovsky (2023). Consideration of the risks, challenges and rewards thus becomes crucial in the ethical deployment of LLMs.
@@ 5 Conclusion
6 Limitations
We acknowledge the limited number of LLMs tested in our work. Though there are many existing models, we opt for the 2 most recognizable representatives of proprietary and open-source models at the time of writing. We encourage other researchers and Machine Learning practitioners to investigate other models of different sizes from alternative platforms.
@@ 6 Limitations
There also exist temporal and geographical constraints. GPT-3.5-Turbo’s cutoff date of their training materials is September 2021; Llama 3 is released in early 2024 Meta (2024). The U.S statistics are available for the years 2022 and 2023. Thus, the LLMs’ knowledge cutoff may be affected after updates. The analysis in our paper is restricted to U.S-based names and statistics. Future studies could expand cross-cultural/national settings to investigate differences in trends.
Finally, we acknowledge that there are multiple ways LLMs could be applied to employment recommendation in practice. Though our work focuses only a number of specific use cases to reveal bias, our findings serves as a cautionary tale on bias for practitioners who desire to utilize LLMs for their applications. A growing body of literature has examined methods to mitigate bias in Machine Learning system Zhou et al. (2023); Zhang et al. (2024). We encourage researchers to peruse these works to reduce bias.
7 Ethics
This work carries minor risks; it identifies challenges with using LLMs in employment decision pipelines which hopefully reduces (rather than exacerbates) such potential uses. It focuses on English only, and biases from a very U.S. perspective, amplifying the exposure of that language/culture. This project did not include data annotation, and only used freely available datasets consistent with their intended uses. AI assistants were not used during this research project or the preparation of this manuscript.
@@ A.2 List of Names used in this Work
    Hispanc Females: Alejandra, Altagracia, Aracelis, Belkis, Denisse, Estefania, Flor, Gisselle, Grisel, Heidy, Ivelisse, Jackeline, Jessenia, Lazara, Lisandra, Luz, Marianela, Maribel, Maricela, Mariela, Marisela, Marisol, Mayra, Migdalia, Niurka, Yaritza, Yesenia, Yessenia, Zoila, Zulma
@@ A.3 LLM Configuration
A.4 BiasinBios Dataset
The BiasinBios dataset, proposed by De-Arteaga et al. (2019), contains English biographies created by the Common Crawl for 28 occupations. For each occupation, there exists a marker that delineates whether the gender of the original owner of the biography. The original biographies have various lengths with a long-tail distribution. Thus, we limit our selections to passages that consist between 80 (the 75% percentile) to 120 words to allow the biographies sufficient space to contain relevant details. We first use GPT-4o (version gpt-4o-2024-05-13) with the prompt template in 9(a) to replace all references to the original personal name with the string "{name}". Then, we use the template in 9(b) to further replace gender-specific pronounces with their gender-neutral counterparts. Finally, we manually go through all 560 rewritten biographies to ensure gender-neutrality while still adhere to relevant details in the original. 10(a) shows a sample data in its original form, and 10(b) shows its rewritten gender-neutral version.
@@ A.4 BiasinBios Dataset
(a) Sample output when original name references are removed.
(b) Sample output rewritten for gender-neutrality.
A.5 Bias Mitigation Strategies
Bias mitigation techniques have garnered much interest in the research community. As our work reveals the potential LLM-propagated inequality in the allocation of employment due to first name preference, the discussion to reduce this bias becomes even more important.
Name-blind Recruitment
@@ A.5 Bias Mitigation Strategies
Bias-aware Finetuning and Prompt Engineering
The first approach involves modifying the LLMs directly to encourage fair behaviors Garimella et al. (2022); Lin et al. (2024). The latter involves modifying the prompt used to interact with the model to reduce bias Li et al. (2024); Dong et al. (2024). These methods could be combined to target bias reduction at multiple checkpoints of LLM-deployment.
Post-hoc Processing

### v2

@@ 1 Introduction
Recently, Large Language Models (LLMs) have become the leading architecture for many tasks in Natural Language Processing (NLP) Kojima et al. (2022); Zhou et al. (2022); Chang et al. (2024). Despite their class-leading performance, LLMs have been shown to propagate and amplify different forms of bias in numerous domains Wan et al. (2023); Gupta et al. (2023); Poulain et al. (2024); Salinas et al. (2023); Li et al. (2024b), similar to how more traditional predictive machine learning-based models replicate and exacerbate social biases Mehrabi et al. (2021).
@@
    Candidates with White-aligned names are preferred by GPT-3.5-Turbo and Llama 3 over other groups in 50% to 95% of 40 occupations, depending on the setting and model.
@@
    Even when candidates possess identical qualifications as reflected in biographies, the average salary offered by these LLMs to candidates with female names may still differ up to 1.8% compared to their male counterpart’s. This discrepancy reaches up to 5% when comparing candidates from intersectional groups.
@@
    Biases exhibited by LLMs partially mirror real-world trends in the United States (U.S) labor force at coarse-grain levels. However, intersectional analysis reveals nuanced discrepancies that favor certain minority groups while punishing others, albeit inconsistently.
Our work builds directly on that of Haim et al. (2024) and of An et al. (2024). Haim et al. (2024) prompted LLMs to provide assistance for 40 Black and White named individuals across topics related to sports, public office, purchasing etc., finding that Black female names received the worst outcomes. An et al. (2024) prompted LLMs to write emails to accept or reject job candidates with stereotypically White, Black or Hispanic names (across two genders), and investigated whether those emails chose to accept or reject the candidates. Their work found that acceptance rates for the latter 2 groups tend to be lower than the former, even when degrees of education and qualification level were consistently stated across candidates. Our work augments these findings by: 1) exploring alternative hiring-related tasks, including salary prediction with full, natural biographies—similar to “résumé studies” in sociology—, and 2) by connecting LLM behaviors to real-world labor data to reveal intersectional bias with respect to a range of occupations.
@@ 2 Hiring Recommendation
In this paper, we study two types of recommendations that LLMs could conceivably be applied. The first, discussed in this section, is hiring recommendations: given an occupation and a list of names of potential candidates, do LLMs exhibit any racial or gender preferences for selecting who to give a job to ? The second type is salary recommendation (§ 3): given a candidate name and (potentially) a biography for that candidate, what salary is recommended for them ?
@@ 2.1 Curation of Names
We limit our study to first (given) names, drawn from two datasets: Rosenman et al. (2023) and Tzioumis (2018). The former contains 136,000 first names compiled form voter registration files while the latter contains 4,250 first names compiled mortgage information in the United States. Both datasets supply the conditional probabilities of each name’s association with 5 racial/ethnic groups: White, Black, Hispanic, Asian, Other. We designate a name as male or female based on its probability in the U.S. Social Security Agency’s database, with a 0.5 probability threshold. Omitting the Other category due to its small size, we randomly select 40 names from each gender of the four remaining races, whose conditional probability P(race|name) exceeds at least 0.8. Our final pool consists of 320 first names. See § A.1 for additional details on the curation process.
@@ 2.3 Hiring Recommendation Results
First, LLMs show a strong preference for White-aligned names, particularly favoring White female names over other groups. For gender-stratified hiring, White female names are preferred in more occupations (35 by GPT-3.5, 29 by Llama 3) compared to White male names (30 and 18) (Table 1). For gender-inclusive hiring, White female names are preferred in 28 (70%) and 26 (65%) occupations by GPT-3.5 and Llama 3 (Table 2)
Second, Llama 3 exhibits less bias for White-aligned names compared to GPT-3.5. In Table 1, Asian names are the second most chosen group across occupations, though not significantly so. In contrast, Black names are disproportionately hired as rapper by GPT-3.5, with the addition of singer and social worker by Llama 3. Hispanic names are never the majority for any occupation by GPT-3.5, and only for 5 and 2 occupations among male and female groups by Llama 3. In Table 2, Llama 3 exhibits more distributed preference for non-White names vs. GPT-3.5, though still far from parity.
@@ 3.2.2 Intersectional Analysis
We see that LLM-projected gender salary gaps are still significantly less than U.S data’s on average. The 2022 ACS reports that females make more than males in only 3 of 16 occupations (dietitian, interior designer, paralegal), with the average absolute percentage gap between the median salaries of the 2 genders at 13.03% (Table 9). In contrast, the average gender gaps between LLMs’ recommended mean salaries are all less than 1.01 $\pm$ 0.82% (Table 6). The average MAEs with respect to U.S statistics remain consistent around 12 units for both LLMs with comparable variance.
@@
4 Bias Mitigation Strategies
As our work reveals the potential LLM-propagated inequality in the allocation of employment due to first name preference, the discussion to reduce this bias becomes even more important. In recent years, bias mitigation techniques have garnered much interest in the research community. We discuss three strategies below that could potentially reduce the observed disparity in LLM-powered hiring.
Name-blind Recruitment
Bias-aware Finetuning and Prompt Engineering
The first approach involves modifying the LLMs directly to encourage fair behaviors Garimella et al. (2022); Lin et al. (2024). The latter involves modifying the prompt used to interact with the model to reduce bias Li et al. (2024a); Dong et al. (2024). These methods could be combined to target bias reduction at multiple checkpoints of deployment.
Post-hoc Processing
@@ 5 Discussion
For instance, female names are preferred over male names in gender-inclusive hiring, yet often offered less salary for the same position than their male counterparts. In contrast, Black names are often overlooked in hiring, but are also offered salary higher than average. In comparison, White-aligned names are consistently preferred in both hiring and salary recommendation, with Hispanic names often on the opposite end. We surmise that this observation may be a potential byproduct of the alignment tuning process that many current LLMs undergo Street (2024); Ouyang et al. (2024). Further investigation is warranted to understand the underlying mechanism of this seemingly counterintuitive artifact.
@@
Our results showcase that prompting LLMs to choose one among several candidates arguably magnify the risk of preferential treatment, and thus should be avoided. Though Llama 3 displays larger magnitude of bias than GPT-3.5, its open-source nature lends itself to more mitigation strategies Zhou et al. (2023); Qureshi et al. (2023); Wang and Russakovsky (2023). Consideration of the risks, challenges and rewards becomes crucial in the ethical deployment of LLMs.
@@ 6 Conclusion
Limitations
We acknowledge the limited number of LLMs tested in our work. Though there are many existing models, we opt for the 2 most recognizable representatives of proprietary and open-source models at the time of writing. We encourage researchers and Machine Learning practitioners to investigate other models from alternative platforms.
@@
There also exist temporal and geographical constraints. GPT-3.5-Turbo’s cutoff date of their training materials is September 2021; Llama 3 is released in early 2024 Meta (2024). The U.S statistics are available for the years 2022 and 2023. Thus, the LLMs’ knowledge cutoff may be affected after updates. The analysis in our paper is restricted to U.S-based names and statistics. It is possible that some of the observed disparity in outcomes by LLMs correlate with the popularity of certain names in the training data. Future studies could expand cross-cultural/national settings to investigate differences in trends.
Finally, we acknowledge that there are multiple ways LLMs could be applied to employment recommendation in practice. Though our work focuses only a number of specific use cases to reveal bias, our findings serves as a cautionary tale on bias for practitioners who desire to utilize LLMs for their applications. We encourage researchers to peruse the growing body of literature on bias mitigation in Machine Learning in their use cases Zhou et al. (2023); Zhang et al. (2024).
Ethics
This work carries minor risks; it identifies challenges with using LLMs in employment decision pipelines which hopefully reduces (rather than exacerbates) such potential uses. It focuses on English only, and biases from a very U.S. perspective, amplifying the exposure of that language/culture. This project did not include data annotation, and only used freely available datasets consistent with their intended uses.
Acknowledgement
This work is funded by the University of Maryland’s Institute for Trustworthy AI in Law & Society (TRAILS). We would like to thank Haozhe An and Rachel Rudinger at the University of Maryland for their support during the conceptualization of this paper. We thank the service of ACL ARR reviewers, area chairs and the editors of the EMNLP conference for our paper’s publication.
@@ A.2 List of Names used in this Work
    Hispanc Females: Alejandra, Altagracia, Aracelis, Belkis, Denisse, Estefania, Flor, Gisselle, Grisel, Heidy, Ivelisse, Jackeline, Jessenia, Lazara, Lisandra, Luz, Marianela, Maribel, Maricela, Mariela, Marisela, Marisol, Mayra, Migdalia, Niurka, Noelia, Odalys, Rocio, Xiomara, Yadira, Yahaira, Yajaira, Yamile, Yanet, Yanira, Yaritza, Yesenia, Yessenia, Zoila, Zulma
@@ A.3 LLM Configuration
A.4 BiasinBios Dataset
The BiasinBios dataset, proposed by De-Arteaga et al. (2019), contains English biographies created by the Common Crawl for 28 occupations. For each occupation, there exists a marker that delineates whether the gender of the original owner of the biography. The original biographies have various lengths with a long-tail distribution. Thus, we limit our selections to passages that consist between 80 (the 75% percentile) to 120 words to allow the biographies sufficient space to contain relevant details. We first use GPT-4o (version gpt-4o-2024-05-13) with the prompt template in 10(a) to replace all references to the original personal name with the string "{name}". Then, we use the template in 10(b) to further replace gender-specific pronounces with their gender-neutral counterparts. Finally, we manually go through all 560 rewritten biographies to ensure gender-neutrality while still adhere to relevant details in the original. 9(a) shows a sample data in its original form, and 9(b) shows its rewritten gender-neutral version.
@@ A.4 BiasinBios Dataset
(a) Sample output when original name references are removed.
(b) Sample output rewritten for gender-neutrality.